# Characterization of a quantum bus between two driven qubits

### v1

@@ IV Implementation of a two-qubit gate
Figure 3: Evolution of the qubit expectation values for an initial state $|10\rangle$ for thermal excitation amplitudes (a) $n_{\mathrm{th},h}=0$ and (b) $n_{\mathrm{th},h}=2$. Entanglement between the different subsystems as a function of time for (c) $n_{\mathrm{th},h}=0$ and (d) $n_{\mathrm{th},h}=2$. The parameters used for all plots are $\Delta=0$, $\Delta_{R}=5\cdot 10^{-2}\omega_{h}$, $g=5\cdot 10^{-3}\omega_{h}$, $\gamma_{q}=10^{-6}\omega_{h}$ and $\gamma_{h}=10^{-6}\omega_{h}$.
In the previous section, we characterize the coupling between the qubit and the harmonic oscillator, highlighting the benefits of using a driven qubit for state readout. In the following, we consider two driven qubits coupled to the same harmonic oscillator, as illustrated in Fig. 1(b). We study the implementation of a two-qubit gate, where the qubit-qubit interaction is mediated by the harmonic oscillator. The Hamiltonian for a two-qubit system can be obtained by a direct generalization of Hamiltonian (3)
@@
Within the rotating wave approximation, the coupling between the qubit and the harmonic oscillator gives rise to an effective qubit-qubit interaction [36, 19, 24], see Appendix B. The interaction time between the qubits can be controlled by tuning their drive amplitudes $\Omega_{R,j}(t)$ close to resonance with the harmonic oscillator. Assuming equal drive amplitudes and coupling strengths for both qubits, the unitary evolution of the system for an interaction time of $t_{\mathrm{int}}=\frac{\pi|\Delta_{R}|}{4g^{2}}$ corresponds to a $\sqrt{i\mathrm{SWAP}}$-gate operation, up to one-qubit transformations,
@@
where the upper and lower sign correspond to a positive and negative Rabi detuning. The $\sqrt{i\mathrm{SWAP}}$-gate generates entanglement between the qubits if they are in opposite logical states. For instance, it maps the separable state $|10\rangle$ into the maximally entangled state $(|10\rangle-i|01\rangle)/\sqrt{2}$. In the following, we study how the qubit-harmonic oscillator coupling enables the qubit-qubit entanglement generation, and analyze the gate fidelity to determine a suitable range of the system parameters for a practical implementation.
@@ IV.1 Qubit entanglement
Universal quantum computation requires the capacity to generate any pure state from an arbitrary initial state. Typically, qubits are initialized in a separable state, which can be manipulated into an entangled state through a sequence of quantum gates. In this section, we analyze the capacity to generate a fully entangled two-qubit state. The qubit-qubit coupling is mediated by a harmonic oscillator, which may become entangled with the qubits during the process. The main goal of this section is to determine whether the qubit-harmonic oscillator system results in qubit-qubit entanglement, or if these are independent processes, and to determine how thermal effects hinder the generation of a fully entangled state. Entanglement in pure bipartite systems is typically measured by the von Neumann entropy of the reduced density matrices. However, for mixed or multipartite systems, the von Neumann entropy is not always a suitable measure, as a classically correlated mixed state can have nonzero von Neumann entropy even though it is not entangled. Several alternative measures of entropy exist for mixed systems [37, 38]. Here, we use logarithmic negativity [39] as a measure of entanglement. Given two subsystems, $A$ and $B$, logarithmic negativity is defined as
@@
where $\hat{\rho}^{\mathrm{T}_{A}}$ is the partial transpose with respect to subsystem $A$ and $||\hat{X}||=\mathrm{Tr}\sqrt{\hat{X}^{\dagger}\hat{X}}$ is the trace norm. Unlike the entropy of entanglement, logarithmic negativity is entanglement monotone [40] both for pure and mixed states. For 2-level systems such as qubits, logarithmic negativity is normalized such that $E_{\mathcal{N}}=1$ for a maximally entangled state.
We solve the quantum master equation (7) for the system Hamiltonian (10[33]. To compute the entanglement between the different subsystems, we perform partial traces over $\hat{\rho}(t)$ to obtain the relevant reduced density matrix. In Fig. 3(a), we show the evolution of the two-qubit states for two consecutive applications ($0\leq t\leq 2t_{\mathrm{int}}$) of the $\sqrt{-i\mathrm{SWAP}}$ gate, with initial state $|10\rangle$ and thermal occupation $n_{\mathrm{th},h}=0$. Ideally, the system should evolve into the $(|10\rangle-i|01\rangle)/\sqrt{2}$ state at $t=t_{\mathrm{int}}$, and to the $-i|01\rangle$ state at $t=2t_{\mathrm{int}}$, swapping the states of the qubits up to a phase factor. On top of the swapping of the qubit occupation numbers $\sigma_{x,i}$, there are small oscillations with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$ and an amplitude of the order of $2g^{2}/\Delta_{R}^{2}$. The cross-correlation $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ oscillates close to -1, since the qubits ideally remain in opposite states. The small oscillations arise due to deviation of the parameters from the dispersive regime, i.e., by retaining higher-order terms in the $g\ll|\Delta_{R}|$ expansion carried out in Appendix B. In the plot, we have chosen values of $g$ and $\Delta_{R}$ satisfying $(\Delta_{R}+4g^{2}/\Delta_{R})t_{\mathrm{int}}\in 2\pi\mathbb{N}$, so that $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ is minimal at $t=t_{\mathrm{int}}$, thus maximizing the gate fidelity $\mathcal{F}$. Nonetheless, the crossing point of the occupation numbers $\langle\sigma_{x}^{1}\rangle=\langle\sigma_{x}^{2}\rangle=0$ occurs at a time slightly higher than $t=t_{\mathrm{int}}$. A finer tuning of $\Delta_{R}$ and $g$ could simultaneously optimize the qubit expectation values to enhance the fidelity.
@@
In Fig. 3(c-d) we show the entanglement, quantified by the logarithmic negativity (12), between the different subsystems, namely the harmonic oscillator ($h$), the subsystem formed by the two qubits ($q$), and each qubit ($q_{1}$ and $q_{2}$). The system begins in a separable state and evolves into a state with maximally entangled qubits at $t=t_{\mathrm{int}}$ (purple line). The harmonic oscillator-qubit entanglement $E_{h,q}$ (yellow line) oscillates periodically with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$, while the entanglement of the harmonic oscillator with each of the individual qubits oscillates in antiphase. The amplitudes of the oscillations of $E_{h,q_{1}}$ (blue line) and $E_{h,q_{2}}$ (red line) are modulated by their respective qubit’s state. Since the qubits are in near resonance with the harmonic oscillator, one may simplify the coupling term in Eq. (10) by applying the rotating wave approximation. Within the rotating wave approximation, doubly rotating terms are disregarded, so that the remaining terms describe energy quanta exchange between the qubit and the harmonic oscillator $\hat{\sigma}_{-,x}^{j}\hat{a}^{\dagger}$ and $\hat{\sigma}_{+,x}^{j}\hat{a})$. Here, $\hat{\sigma}_{\pm,x}$ are the rising and lowering operators in the dressed qubit’s basis. Since at $T=0$ the harmonic oscillator is initialized in the ground state, at $t=0$ it may only couple to the first qubit, which starts in the excited state $|1\rangle$. The situation is reversed at $t=2t_{\mathrm{int}}$.
At a nonvanishing temperature, Fig. 3(d), in agreement with the results of Fig. 3(b), the oscillations of the entanglement get suppressed as time progresses. However, close to $t=2t_{\mathrm{int}}$, higher order harmonics become relevant. While entanglement is not a conserved quantity, the general trend in Figs. 3(c-d) is the pumping of the $h-q_{1}$ entanglement into the $q_{1}-q_{2}$ entanglement for $t\lesssim t_{\mathrm{int}}$, and the pumping of $E_{q_{1},q_{2}}$ to $h-q_{2}$ entanglement for $t_{\mathrm{int}}\lesssim t\lesssim 2t_{\mathrm{int}}$. Therefore, the suppression of the $h-q_{j}$ entanglement at high temperatures lowers the ability of the quantum gate to generate a maximally entangled $q_{1}-q_{2}$ state.
@@ IV.2 Fidelity of the two-qubit gate
Figure 4: (a) Fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate as a function of the driving field detuning and the Rabi detuning. The parameters used are $n_{\mathrm{th},h}=0$, $g=5\cdot 10^{-3}\omega_{h}$, $\gamma_{q}=10^{-6}\omega_{h}$ and $\gamma_{h}=10^{-6}\omega_{h}$. (b) Fidelity of the quantum gate as a function of the qubit and harmonic oscillator dampings. The parameters used are $n_{\mathrm{th},h}=0$, $\Delta=0$, $\Delta_{R}=5\cdot 10^{-2}\omega_{h}$ and $g=5\cdot 10^{-3}\omega_{h}$. (c) Depiction of the dressed qubit transition frequency for a nonvanishing dressing rise time, and (d) fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate as a function of the rise time. The parameters used are $n_{\mathrm{th},h}=0$, $\Delta=0$, $\gamma_{q}=10^{-6}\omega_{h}$ and $\gamma_{h}=10^{-6}\omega_{h}$.
In the previous section we study how temperature affects the qubit-qubit entanglement generation, which is crucial for universal quantum computation. The thermal occupation of the system is determined by the temperature of its environment. As a rule of thumb, higher temperatures are detrimental to coherence. The system parameters, such as characteristic frequencies and coupling constants, are also critical in achieving high-fidelity quantum computing. Typically, these parameters are fixed by external electric or magnetic fields, materials choices, or the physical realization of the quantum register. In the following, we study how different parameters of the model affect the fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate.
Fidelity $\mathcal{F}$ is a measure of how closely a quantum gate’s implementation matches the ideal behavior ($\mathcal{F}=1$). For pure states, it is given by the inner product of the real and ideal final states, averaged over all initial pure states. However, thermal fluctuations and damping drive the system into a mixed state, so the definition of the gate fidelity takes a more general form:
@@
Fidelity is non-monotonous in $\Delta_{R}$. On the one hand, the rotating wave approximation necessary for building the harmonic oscillator-mediated $\sqrt{i\mathrm{SWAP}}$-gate [see Appendix A] requires a small Rabi detuning $|\Delta_{R}|\ll\omega_{h}$. On the other hand, the dispersive regime is valid for $g\ll\Delta_{R}$. Focusing on $\Delta=0$ (blue line), the rotating wave approximation is well satisfied for Rabi detunings as large as $|\Delta_{R}|\approx 0.1\omega_{h}$, but the fidelity has a significant drop for $|\Delta_{R}|/g<10$ due to deviations from the dispersive regime. In the resonant regime $\Delta_{R}\ll g$, the harmonic oscillator may adiabatically be eliminated assuming that the harmonic oscillator begins in its ground state. The effective qubit-qubit interaction may also be used to build a $\sqrt{i\mathrm{SWAP}}$-gate. However, the qubits are more prone to decoherence in resonant regime. Regarding the driving field detuning, a nonvanishing $\Delta$ changes the qubits’ eigenbasis. Therefore, a high-fidelity gate implementation requires that $\Delta\ll\Omega_{R},\omega_{h}$. Based on the results of Fig. 4(a), an appropriate range of values for the detunings is $\Delta<0.01\omega_{h}$ and $10g<|\Delta_{R}|\lesssim 0.1\omega_{h}$. For a given $\Delta$, $|\Delta_{R}|$ may be calibrated to maximize $\mathcal{F}$.
In Fig. 4(b) we study the dependence of the gate fidelity on the damping rates. The damping rates quantify the coupling of the system to the environment, so higher damping rates will always amount to lower $\mathcal{F}$. Although the dressed qubit and harmonic oscillator characteristic frequencies are approximately the same $\Omega_{R}\sim\omega_{h}$, the qubit damping rate has a more critical impact on $\mathcal{F}$. The fidelity decreases significantly for values higher than $\gamma_{q}\gtrsim 10^{-5}\omega_{h}$, while the harmonic oscillator rate can be as high as $\gamma_{h}=10^{-3}\omega_{h}$ to account for the same fidelity loss. Taking into account that the bare qubit transition frequency is higher than the dressed qubit one $\omega_{q}\gg\Omega_{R},\omega_{h}$, the qubit quality factor $Q_{q}=\omega_{q}/\gamma_{q}\gg 10^{5}$ must be remarkably high to avoid fidelity loss due to the coupling to the environment.
@@
as depicted in Fig. 4(c). In Fig. 4(d) we show how the gate fidelity depends on the rise time. Based on the previous analysis of the system parameters, we select values of the detunings and coupling constant suitable for quantum computation. Modern quantum computers achieve two-qubit gate fidelities of approximately $\mathcal{F}\sim 99\%$ [41, 42, 43]. A minimum fidelity of $\mathcal{F}\sim 99.99\%$ [44, 45], combined with quantum error correction protocols, is expected for practical quantum computing. The general trend shows that gate fidelity decreases with slower rise times, but there is also an oscillatory behavior that may be exploited to enhance $\mathcal{F}$. Given the practical limitations on the rise time, one may calibrate the rise time to maximize the fidelity.
V Practical realization
Most present-day quantum computer realizations rely on qubits coupled via electromagnetic resonators or with other qubits connected to each other electromagnetically [30]. Such interqubit coupling requires that the qubit states are discernible based on their different charge susceptibilities. Those qubits are inherently susceptible to offset charge fluctuations, even if the effective coupling to those fluctuations can be minimized as done, for example, in superconducting transmon qubits [46]. The longest single-qubit coherence times are therefore obtained with qubits having no coupling to charge, such as spin qubits [47]. Then the problem is to find a suitable mediator for coupling such qubits while maintaining their good coherence. One possible choice is based on utilizing mechanical vibrations in a mechanical resonator, coupling to the qubit states via any mechanism where the resonator’s motion shifts the (Zeeman) spin splitting, such as the qubit strain-induced effects [48], or a motion-modulated external magnetic field. The latter mechanism hence requires a spatial magnetic field gradient, and the coupling strength is proportional to the amount of change in the magnetic field within the zero-point vibration amplitude.

### v2

@@ IV Implementation of a two-qubit gate
Figure 3: Evolution of the qubit expectation values for an initial state $|10\rangle$ for thermal excitation amplitudes (a) $n_{\mathrm{th},h}=0$ and (b) $n_{\mathrm{th},h}=2$. Entanglement between the different subsystems as a function of time for (c) $n_{\mathrm{th},h}=0$ and (d) $n_{\mathrm{th},h}=2$. All plots use the parameters $\Delta=0$, $\Delta_{R}=5\cdot 10^{-2}\omega_{h}$, $g=5\cdot 10^{-3}\omega_{h}$, $\gamma_{q}=10^{-6}\omega_{h}$ and $\gamma_{h}=10^{-6}\omega_{h}$.
In the previous section, we characterise the coupling between the qubit and the harmonic oscillator, highlighting the benefits of using a driven qubit for state readout. In the following, we consider two driven qubits coupled to the same harmonic oscillator, as illustrated in Fig. 1(b). We study the implementation of a two-qubit gate, where the qubit-qubit interaction is mediated by the harmonic oscillator. The Hamiltonian for a two-qubit system can be obtained by a direct generalisation of Hamiltonian (3)
@@
Within the rotating wave approximation, the coupling between the qubit and the harmonic oscillator gives rise to an effective qubit-qubit interaction [38, 19, 24], see Supplementary note C [29]. The interaction time between the qubits can be controlled by tuning their drive amplitudes $\Omega_{R,j}(t)$ close to resonance with the harmonic oscillator. Assuming equal drive amplitudes and coupling strengths for both qubits, the unitary evolution of the system for an interaction time of $t_{\mathrm{int}}=\frac{\pi|\Delta_{R}|}{4g^{2}}$ corresponds to a $\sqrt{i\mathrm{SWAP}}$-gate operation, up to one-qubit transformations,
@@
where the upper and lower signs correspond to a positive and negative Rabi detuning. The $\sqrt{i\mathrm{SWAP}}$-gate generates entanglement between the qubits if they are in opposite logical states. For instance, it maps the separable state $|10\rangle$ into the maximally entangled state $(|10\rangle+i|01\rangle)/\sqrt{2}$. In Supplementary note D [29] we show how to create Bell states in driven qubits. In the following, we study how the qubit-harmonic oscillator coupling enables the qubit-qubit entanglement generation, and analyze the gate fidelity to determine a suitable range of the system parameters for a practical implementation.
@@ IV.1 Qubit entanglement
Universal quantum computation requires the capacity to generate any pure state from an arbitrary initial state. Typically, qubits are initialised in a separable state, which can be manipulated into an entangled state through a sequence of quantum gates. In this section, we analyze the capacity to generate a fully entangled two-qubit state. The qubit-qubit coupling is mediated by a harmonic oscillator, which may become entangled with the qubits during the process. The main goal of this section is to determine whether the qubit-harmonic oscillator system results in qubit-qubit entanglement, or if these are independent processes, and to determine how thermal effects hinder the generation of a fully entangled state. Entanglement in pure bipartite systems is typically measured by the von Neumann entropy of the reduced density matrices. However, for mixed or multipartite systems, the von Neumann entropy is not always a suitable measure, as a classically correlated mixed state can have nonzero von Neumann entropy even though it is not entangled. Several alternative measures of entropy exist for mixed systems [39, 40]. Here, we use logarithmic negativity [41, 42] as a measure of entanglement. Given two subsystems, $A$ and $B$, logarithmic negativity is defined as
@@
where $\hat{\rho}^{\mathrm{T}_{A}}$ is the partial transpose with respect to subsystem $A$ and $||\hat{X}||=\mathrm{Tr}\sqrt{\hat{X}^{\dagger}\hat{X}}$ is the trace norm. Unlike the entropy of entanglement, logarithmic negativity is entanglement monotone [43, 42] both for pure and mixed states. For 2-level systems such as qubits, logarithmic negativity is normalised such that $E_{\mathcal{N}}=1$ for a maximally entangled state.
We solve the quantum master equation (7) for the system Hamiltonian (10[35]. To compute the entanglement between the different subsystems, we perform partial traces over $\hat{\rho}(t)$ to obtain the relevant reduced density matrix. In Fig. 3(a), we show the evolution of the two-qubit states for two consecutive applications ($0\leq t\leq 2t_{\mathrm{int}}$) of the $\sqrt{-i\mathrm{SWAP}}$ gate, with initial state $|10\rangle$ and thermal occupation $n_{\mathrm{th},h}=0$. Ideally, the system should evolve into the $(|10\rangle-i|01\rangle)/\sqrt{2}$ state at $t=t_{\mathrm{int}}$, and to the $-i|01\rangle$ state at $t=2t_{\mathrm{int}}$, swapping the states of the qubits up to a phase factor. On top of the swapping of the qubit occupation numbers $\sigma_{x,i}$, there are small oscillations with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$ and an amplitude of the order of $2g^{2}/\Delta_{R}^{2}$. The cross-correlation $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ oscillates close to -1, since the qubits ideally remain in opposite states. The small oscillations arise due to deviation of the parameters from the dispersive regime, i.e., by retaining higher-order terms in the $g\ll|\Delta_{R}|$ expansion carried out in Supplementary note C [29]. In the plot, we have chosen values of $g$ and $\Delta_{R}$ satisfying $(\Delta_{R}+4g^{2}/\Delta_{R})t_{\mathrm{int}}\in 2\pi\mathbb{N}$, so that $\langle\sigma_{x}^{1}\sigma_{x}^{2}\rangle$ is minimal at $t=t_{\mathrm{int}}$, thus maximizing the gate fidelity $\mathcal{F}$. Nonetheless, the crossing point of the occupation numbers $\langle\sigma_{x}^{1}\rangle=\langle\sigma_{x}^{2}\rangle=0$ occurs at a time slightly higher than $t=t_{\mathrm{int}}$. A finer tuning of $\Delta_{R}$ and $g$ could simultaneously optimise the qubit expectation values to enhance the fidelity.
@@
In Fig. 3(c-d) we show the entanglement, quantified by the logarithmic negativity (12), between the different subsystems, namely the harmonic oscillator ($h$), the subsystem formed by the two qubits ($q$), and each qubit ($q_{1}$ and $q_{2}$). The system begins in a separable state and evolves into a state with maximally entangled qubits at $t=t_{\mathrm{int}}$ (purple line). The harmonic oscillator-qubit entanglement $E_{h,q}$ (yellow line) oscillates periodically with frequency $\Delta_{R}+4g^{2}/\Delta_{R}$, while the entanglement of the harmonic oscillator with each of the individual qubits oscillates in antiphase. The amplitudes of the oscillations of $E_{h,q_{1}}$ (blue line) and $E_{h,q_{2}}$ (red line) are modulated by their respective qubit’s state. Since the qubits are in near resonance with the harmonic oscillator, one may simplify the coupling term in Eq. (10) by applying the rotating wave approximation. Within the rotating wave approximation, doubly rotating terms are disregarded, so that the remaining terms describe energy quanta exchange between the qubit and the harmonic oscillator ($\hat{\sigma}_{-,x}^{j}\hat{a}^{\dagger}$ and $\hat{\sigma}_{+,x}^{j}\hat{a})$. Here, $\hat{\sigma}_{\pm,x}$ are the rising and lowering operators in the dressed qubit’s basis. Since at $T=0$ the harmonic oscillator is initialised in the ground state, at $t=0$ it may only couple to the first qubit, which starts in the excited state $|1\rangle$. The situation is reversed at $t=2t_{\mathrm{int}}$.
At a nonvanishing temperature, Fig. 3(d), in agreement with the results of Fig. 3(b), the oscillations of the entanglement get suppressed as time progresses. However, close to $t=2t_{\mathrm{int}}$, higher order harmonics become relevant. While entanglement is not a conserved quantity, the general trend in Figs. 3(c-d) is the pumping of the $h-q_{1}$ entanglement into the $q_{1}-q_{2}$ entanglement for $t\lesssim t_{\mathrm{int}}$, and the pumping of $E_{q_{1},q_{2}}$ to $h-q_{2}$ entanglement for $t_{\mathrm{int}}\lesssim t\lesssim 2t_{\mathrm{int}}$. Therefore, the suppression of the $h-q_{j}$ entanglements at high temperatures lowers the ability of the quantum gate to generate a maximally entangled $q_{1}-q_{2}$ state.
@@ IV.2 Fidelity of the two-qubit gate
Figure 4: (a) Fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate as a function of the driving field detuning and the Rabi detuning. The parameters used are $n_{\mathrm{th},h}=0$, $g=5\cdot 10^{-3}\omega_{h}$, $\gamma_{q}=10^{-6}\omega_{h}$ and $\gamma_{h}=10^{-6}\omega_{h}$. (b) Fidelity of the quantum gate as a function of the qubit and harmonic oscillator dampings. The parameters used are $n_{\mathrm{th},h}=0$, $\Delta=0$, $\Delta_{R}=5\cdot 10^{-2}\omega_{h}$ and $g=5\cdot 10^{-3}\omega_{h}$. (c) Depiction of the dressed qubit transition frequency for a nonvanishing dressing rise time, and (d) fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate as a function of the rise time. All plots use the parameters $n_{\mathrm{th},h}=0$, $\Delta=0$, $\gamma_{q}=10^{-6}\omega_{h}$ and $\gamma_{h}=10^{-6}\omega_{h}$.
In the previous section we study how temperature affects the qubit-qubit entanglement generation, which is crucial for universal quantum computation. The thermal occupation of the system is determined by the temperature of its environment. As a rule of thumb, higher temperatures are detrimental to coherence. The system parameters, such as characteristic frequencies and coupling constants, are also critical in achieving high-fidelity quantum computing. Typically, these parameters are fixed by external electric or magnetic fields, materials choices, or the physical realisation of the quantum register. In the following, we study how different parameters of the model affect the fidelity of the $\sqrt{i\mathrm{SWAP}}$-gate.
Fidelity $\mathcal{F}$ is a measure of how closely a quantum gate’s implementation matches the ideal behaviour ($\mathcal{F}=1$). For pure states, it is given by the inner product of the real and ideal final states, averaged over all initial pure states. However, thermal fluctuations and damping drive the system into a mixed state, so the definition of the gate fidelity takes a more general form:
@@
Fidelity is non-monotonous in $\Delta_{R}$. On the one hand, the rotating wave approximation necessary for building the harmonic oscillator-mediated $\sqrt{i\mathrm{SWAP}}$-gate [see Supplementary note A [29]], requires a small Rabi detuning $|\Delta_{R}|\ll\omega_{h}$. On the other hand, the dispersive regime is valid for $g\ll\Delta_{R}$. Focusing on $\Delta=0$ (blue line), the rotating wave approximation is well satisfied for Rabi detunings as large as $|\Delta_{R}|\approx 0.1\omega_{h}$, but the fidelity has a significant drop for $|\Delta_{R}|/g<10$ due to deviations from the dispersive regime. Regarding the driving field detuning, a nonvanishing $\Delta$ changes the qubits’ eigenbasis. Therefore, a high-fidelity gate implementation requires that $\Delta\ll\Omega_{R},\omega_{h}$. Based on the results of Fig. 4(a), an appropriate range of values for the detunings is $\Delta<0.01\omega_{h}$ and $10g<|\Delta_{R}|\lesssim 0.1\omega_{h}$. For a given $\Delta$, $|\Delta_{R}|$ may be calibrated to maximise $\mathcal{F}$.
In Fig. 4(b) we study the dependence of the gate fidelity on the damping rates. The damping rates quantify the coupling of the system to the environment, so higher damping rates will always amount to lower $\mathcal{F}$. Although the dressed qubit and harmonic oscillator characteristic frequencies are approximately the same $\Omega_{R}\sim\omega_{h}$, the qubit damping rate has a more critical impact on $\mathcal{F}$, since it directly acts on the qubit degrees of freedoms. The inverse qubit damping rate defines the qubit decoherence time, so the damping rate should satisfy $\gamma_{q}t_{\mathrm{int}}\ll 1$ to ensure that the qubit remains coherent throughout the application of the quantum gate. Numerical results show that the fidelity decreases significantly for values exceeding $\gamma_{q}\gtrsim 10^{-2}g^{2}/\Delta_{R}$, whereas the harmonic oscillator rate can be as high as $\gamma_{h}=2g^{2}/\Delta_{R}$ while producing the same fidelity loss.
@@
as depicted in Fig. 4(c). In Fig. 4(d) we show how the gate fidelity depends on the rise time. Based on the previous analysis of the system parameters, we select values of the detunings and coupling constant suitable for quantum computation. Modern quantum computers achieve two-qubit gate fidelities of approximately $\mathcal{F}\sim 99\%$ [44, 45, 46]. A minimum fidelity of $\mathcal{F}\sim 99.99\%$ [47, 48], combined with quantum error correction protocols, is expected for practical quantum computing. The general trend shows that gate fidelity decreases with slower rise times, but there is also an oscillatory behaviour that may be exploited to enhance $\mathcal{F}$. Given the practical limitations on the rise time, one may calibrate the rise time to maximise the fidelity.
V Practical realisation
Most present-day quantum computer realisations rely on qubits coupled via electromagnetic resonators or with other qubits connected to each other electromagnetically [32]. Such interqubit coupling requires that the qubit states are discernible based on their different charge susceptibilities. Those qubits are inherently susceptible to offset charge fluctuations, even if the effective coupling to those fluctuations can be minimised as done, for example, in superconducting transmon qubits [49]. The longest single-qubit coherence times are therefore obtained with qubits having no coupling to charge, such as spin qubits [50]. Then, the problem is to find a suitable mediator for coupling such qubits while maintaining their good coherence. One possible choice is based on utilizing mechanical vibrations in a mechanical resonator, coupling to the qubit states via any mechanism where the resonator’s motion shifts the (Zeeman) spin splitting, such as the qubit strain-induced effects [51], or a motion-modulated external magnetic field. The latter mechanism hence requires a spatial magnetic field gradient, and the coupling strength is proportional to the amount of change in the magnetic field within the zero-point vibration amplitude.